# Finite Sample Analysis of Distributional TD Learning with Linear Function Approximation

### v1

@@ 6 Conclusions
In this paper, we have bridged a critical theoretical gap in distributional reinforcement learning by establishing the non-asymptotic sample complexity of distributional TD learning with linear function approximation. Specifically, we have proposed Linear-CTD, derived via solving the linear-categorical projected Bellman equation with linear stochastic approximation. By carefully analyzing the Bellman equation and using the exponential stability arguments, we have shown tight sample complexity bounds for the proposed algorithm in both generative model setting and Markovian setting. Our finite-sample rates match the state-of-the-art sample complexity bounds for the classic TD learning. These theoretical findings demonstrate that learning the full return distribution under linear function approximation can be statistically as easy as classic TD learning for value function estimation. Finally, we noted that it is possible to improve the convergence rates by applying variance-reduction techniques or using polynomial-decaying step sizes, which we leave as a future work.
@@ Appendix A Kronecker Product
In this section, we will introduce some properties of Kronecker product used in our paper. See [Zhang and Ding, 2013] for a detailed treatment of Kronecker product.
@@
See [Basic properties and Theorem 3 Zhang and Ding, 2013]. ∎
@@
For any matrices ${\bm{B}}_{1},{\bm{B}}_{2},{\bm{B}}_{3}$ such that ${\bm{B}}_{1}{\bm{B}}_{2}{\bm{B}}_{3}$ can be defined, it holds that $\mathsf{vec}\left({{\bm{B}}_{1}{\bm{B}}_{2}{\bm{B}}_{3}}\right)=\left({{\bm{B}}_{3}^{\intercal}\otimes{\bm{B}}_{1}}\right)\mathsf{vec}\left({{\bm{B}}_{2}}\right)$.
@@
See [Lemma 4.3.1 Horn and Johnson, 1994]. ∎
@@
For any matrices ${\bm{B}}_{1}$ and ${\bm{B}}_{2}$, it holds that $\left\|{{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right\|=\left\|{{\bm{B}}_{1}}\right\|\left\|{{\bm{B}}_{2}}\right\|$, $\left({{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right)^{\intercal}={\bm{B}}_{1}^{\intercal}\otimes{\bm{B}}_{2}^{\intercal}$. Furthermore, if ${\bm{B}}_{1}$ and ${\bm{B}}_{2}$ are invertible/orthogonal/diagonal/symmetric/normal, ${\bm{B}}_{1}\otimes{\bm{B}}_{2}$ is also invertible/orthogonal/diagonal/symmetric/normal and $\left({{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right)^{-1}={\bm{B}}_{1}^{-1}\otimes{\bm{B}}_{2}^{-1}$.
@@
See [Basic properties, Theorem 5 and Theorem 7 Zhang and Ding, 2013]. ∎
@@
Consider the spectral decomposition ${\bm{B}}_{i}={\bm{Q}}_{i}{\bm{D}}_{i}{\bm{Q}}_{i}^{\intercal}$, for any $i\in[4]$, by Lemma A.1 and Lemma A.3, we have
$\left({{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right)=\left({{\bm{Q}}_{1}\otimes{\bm{Q}}_{2}}\right)\left({{\bm{D}}_{1}\otimes{\bm{D}}_{2}}\right)\left({{\bm{Q}}_{1}\otimes{\bm{Q}}_{2}}\right)^{\intercal}$
@@
$\left({{\bm{B}}_{3}\otimes{\bm{B}}_{4}}\right)=\left({{\bm{Q}}_{3}\otimes{\bm{Q}}_{4}}\right)\left({{\bm{D}}_{3}\otimes{\bm{D}}_{4}}\right)\left({{\bm{Q}}_{3}\otimes{\bm{Q}}_{4}}\right)^{\intercal}$
@@
${\bm{B}}_{2}\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)^{\intercal}={\bm{B}}_{2}\otimes{\bm{v}}^{\intercal},$
$\left({{\bm{I}}_{K}\otimes{\bm{u}}}\right){\bm{B}}_{3}\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)^{\intercal}={\bm{B}}_{3}\otimes\left({{\bm{u}}{\bm{v}}^{\intercal}}\right).$
@@
${\bm{B}}_{2}\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)^{\intercal}=$ $\left[{\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)\otimes{\bm{B}}^{\intercal}_{2}}\right]^{\intercal}=\left[{{\bm{B}}^{\intercal}_{2}\otimes{\bm{v}}}\right]^{\intercal}={\bm{B}}_{2}\otimes{\bm{v}}^{\intercal}.$
@@
$\left({{\bm{I}}_{K}\otimes{\bm{u}}}\right){\bm{B}}_{3}\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)^{\intercal}=$ $\left({{\bm{B}}_{3}\otimes{\bm{u}}}\right)\otimes{\bm{v}}^{\intercal}={\bm{B}}_{3}\otimes\left({{\bm{u}}\otimes{\bm{v}}^{\intercal}}\right)={\bm{B}}_{3}\otimes\left({{\bm{u}}{\bm{v}}^{\intercal}}\right).$
@@
Appendix B Omitted Results and Proofs in Section 2
B.1 Linear Projected Bellman Equation
@@
$=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({{\mathbb{E}}\left[{r_{0}\mid s_{0}=s}\right]+\gamma{\mathbb{E}}\left[{{\bm{\phi}}(s_{1})^{\intercal}\mid s_{0}=s}\right]{\bm{\psi}}}\right)}\right]$
$=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s,s^{\prime}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right]{\bm{\psi}}+{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s,r}\left[{{\bm{\phi}}(s)r}\right],$
@@
$\left({{\bm{\Sigma}}_{{\bm{\phi}}}-\gamma{\mathbb{E}}_{s,s^{\prime}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right]}\right){\bm{\psi}}={\mathbb{E}}_{s,r}\left[{{\bm{\phi}}(s)r}\right].$
B.2 Convergence Results for Linear TD Learning
@@
In Linear-TD, our goal is to find a good estimator $\hat{{\bm{\psi}}}$ such that $\left\|{{\bm{V}}_{\hat{{\bm{\psi}}}}-{\bm{V}}_{{\bm{\psi}}^{\star}}}\right\|_{\mu_{\pi}}=\left\|{\hat{{\bm{\psi}}}-{\bm{\psi}}^{\star}}\right\|_{{\bm{\Sigma}}_{{\bm{\phi}}}}\leq\varepsilon$. Samsonov et al. [2024b] considered the Polyak-Ruppert tail averaging $\bar{{\bm{\psi}}}_{T}:=\left({T/2+1}\right)^{-1}\sum_{t=T/2}^{T}{\bm{\psi}}_{t}$, and showed that in the generative model setting with constant step size $\alpha\simeq(1-\gamma)\lambda_{\min}$,
$T=\widetilde{O}\left({\frac{\left\|{{\bm{\psi}}^{\star}}\right\|^{2}_{{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)^{2}\lambda_{\min}}\left({\frac{1}{\varepsilon^{2}}+\frac{1}{\lambda_{\min}}}\right)}\right)$
is sufficient to guarantee that $\left\|{{\bm{V}}_{\bar{{\bm{\psi}}}_{T}}-{\bm{V}}_{{\bm{\psi}}^{\star}}}\right\|_{\mu_{\pi}}\leq\varepsilon$. They also provided sample complexity bounds when taking the instance-independent (i.e. not dependent on unknown quantity) step size, and in the Markovian setting.
B.3 Categorical Parametrization is an Isometry
Proposition B.1.
The affine space $\left({{\mathscr{P}}^{\mathsf{sign}}_{K},\ell_{2}}\right)$ is isometric with $\left({{\mathbb{R}}^{K},\sqrt{\iota_{K}}\left\|{\cdot}\right\|_{{\bm{C}}^{\intercal}{\bm{C}}}}\right)$, in the sense that, for any $\nu_{{\bm{p}}_{1}},\nu_{{\bm{p}}_{2}}\in{\mathscr{P}}^{\mathsf{sign}}_{K}$, it holds that $\ell_{2}^{2}(\nu_{{\bm{p}}_{1}},\nu_{{\bm{p}}_{2}})=\iota_{K}\left\|{{\bm{p}}_{1}-{\bm{p}}_{2}}\right\|^{2}_{{\bm{C}}^{\intercal}{\bm{C}}}$, where
${\bm{C}}=\begin{bmatrix}1&0&\cdots&0&0\\ 1&1&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 1&1&\cdots&1&0\\ 1&1&\cdots&1&1\end{bmatrix}\in{\mathbb{R}}^{K\times K}.$ (16)
@@ B.3 Categorical Parametrization is an Isometry
$=$ $\iota_{K}\left\|{{\bm{p}}_{1}-{\bm{p}}_{2}}\right\|^{2}_{{\bm{C}}^{\intercal}{\bm{C}}}.$
B.4 Categorical Projection Operator is Orthogonal Projection
Proposition B.2.
[Bellemare et al., 2023, Lemma 9.17] For any $\nu\in{\mathscr{P}}^{\mathsf{sign}}$ and $\nu_{{\bm{p}}}\in{\mathscr{P}}^{\mathsf{sign}}_{K}$, it holds that
@@ B.4 Categorical Projection Operator is Orthogonal Projection
B.5 Categorical Projected Bellman Operator
@@
${\bm{p}}_{\bm{\mathcal{T}}^{\pi}\bm{\eta}}(s)=$ ${\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r_{0})}\right){\bm{p}}_{\bm{\eta}}(s_{1})\Big{|}s_{0}=s}\right]$
@@
This proposition is a special case of Proposition C.3, whose proof can be found in Appendix C.4.
Appendix C Omitted Results and Proofs in Section 3
C.1 Linear-Categorical Parametrization is an Isometry
Proposition C.1.
The affine space $\left({{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K},\ell_{2,\mu_{\pi}}}\right)$ is isometric with $\left({{\mathbb{R}}^{dK},\sqrt{\iota_{K}}\left\|{\cdot}\right\|_{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}}\right)$, in the sense that, for any $\bm{\eta}_{{\bm{w}}_{1}},\bm{\eta}_{{\bm{w}}_{2}}\in{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}$, it holds that $\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{w}}_{1}},\bm{\eta}_{{\bm{w}}_{2}}}\right)=\iota_{K}\left\|{{\bm{w}}_{1}-{\bm{w}}_{2}}\right\|^{2}_{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}$.

### v2

@@ 6 Conclusions
In this paper, we have bridged a critical theoretical gap in distributional reinforcement learning by establishing the non-asymptotic sample complexity of distributional TD learning with linear function approximation. Specifically, we have proposed Linear-CTD, which is derived by solving the linear-categorical projected Bellman equation. By carefully analyzing the Bellman equation and using the exponential stability arguments, we have shown tight sample complexity bounds for the proposed algorithm. Our finite-sample rates match the state-of-the-art sample complexity bounds for conventional TD learning. These theoretical findings demonstrate that learning the full return distribution under linear function approximation can be statistically as easy as conventional TD learning for value function estimation. Our numerical experiments have provided empirical validation of our theoretical results. Finally, we have noted that it would be possible to improve the convergence rates by applying variance-reduction techniques or using polynomial-decaying step sizes, which we leave for future work.
@@ Appendix A Kronecker Product
In this section, we will introduce some properties of Kronecker product used in our paper. See Zhang and Ding [72] for a detailed treatment of Kronecker product.
@@
See Basic properties and Theorem 3 in Zhang and Ding [72]. ∎
@@
For any matrices ${\bm{B}}_{1},{\bm{B}}_{2},{\bm{B}}_{3}$ such that ${\bm{B}}_{1}{\bm{B}}_{2}{\bm{B}}_{3}$ can be defined, it holds that $\mathsf{vec}\left({{\bm{B}}_{1}{\bm{B}}_{2}{\bm{B}}_{3}}\right)=\left({{\bm{B}}_{3}^{\top}\otimes{\bm{B}}_{1}}\right)\mathsf{vec}\left({{\bm{B}}_{2}}\right)$.
@@
See Lemma 4.3.1 in Horn and Johnson [19]. ∎
@@
For any matrices ${\bm{B}}_{1}$ and ${\bm{B}}_{2}$, it holds that $\left\|{{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right\|=\left\|{{\bm{B}}_{1}}\right\|\left\|{{\bm{B}}_{2}}\right\|$, $\left({{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right)^{\top}={\bm{B}}_{1}^{\top}\otimes{\bm{B}}_{2}^{\top}$. Furthermore, if ${\bm{B}}_{1}$ and ${\bm{B}}_{2}$ are invertible/orthogonal/diagonal/symmetric/normal, ${\bm{B}}_{1}\otimes{\bm{B}}_{2}$ is also invertible/orthogonal/diagonal/symmetric/normal and $\left({{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right)^{-1}={\bm{B}}_{1}^{-1}\otimes{\bm{B}}_{2}^{-1}$.
@@
See Basic properties, Theorem 5 and Theorem 7 in Zhang and Ding [72]. ∎
@@
Consider the spectral decomposition ${\bm{B}}_{i}={\bm{Q}}_{i}{\bm{D}}_{i}{\bm{Q}}_{i}^{\top}$, for any $i\in[4]$, by Lemma A.1 and Lemma A.3, we have
$\left({{\bm{B}}_{1}\otimes{\bm{B}}_{2}}\right)=\left({{\bm{Q}}_{1}\otimes{\bm{Q}}_{2}}\right)\left({{\bm{D}}_{1}\otimes{\bm{D}}_{2}}\right)\left({{\bm{Q}}_{1}\otimes{\bm{Q}}_{2}}\right)^{\top}$
@@
$\left({{\bm{B}}_{3}\otimes{\bm{B}}_{4}}\right)=\left({{\bm{Q}}_{3}\otimes{\bm{Q}}_{4}}\right)\left({{\bm{D}}_{3}\otimes{\bm{D}}_{4}}\right)\left({{\bm{Q}}_{3}\otimes{\bm{Q}}_{4}}\right)^{\top}$
@@
${\bm{B}}_{2}\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)^{\top}={\bm{B}}_{2}\otimes{\bm{v}}^{\top},$
$\left({{\bm{I}}_{K}\otimes{\bm{u}}}\right){\bm{B}}_{3}\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)^{\top}={\bm{B}}_{3}\otimes\left({{\bm{u}}{\bm{v}}^{\top}}\right).$
@@
${\bm{B}}_{2}\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)^{\top}=$ $\left[{\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)\otimes{\bm{B}}^{\top}_{2}}\right]^{\top}=\left[{{\bm{B}}^{\top}_{2}\otimes{\bm{v}}}\right]^{\top}={\bm{B}}_{2}\otimes{\bm{v}}^{\top}.$
@@
$\left({{\bm{I}}_{K}\otimes{\bm{u}}}\right){\bm{B}}_{3}\left({{\bm{I}}_{K}\otimes{\bm{v}}}\right)^{\top}=$ $\left({{\bm{B}}_{3}\otimes{\bm{u}}}\right)\otimes{\bm{v}}^{\top}={\bm{B}}_{3}\otimes\left({{\bm{u}}\otimes{\bm{v}}^{\top}}\right)={\bm{B}}_{3}\otimes\left({{\bm{u}}{\bm{v}}^{\top}}\right).$
@@
Appendix B Related Work
Distributional Reinforcement Learning.
Distributional TD learning was first proposed in Bellemare et al. [2]. Following the distributional perspective in Bellemare et al. [2], Qu et al. [46] proposed a distributional version of the gradient TD learning algorithm, Tang et al. [59] proposed a distributional version of multi-step TD learning, Tang et al. [60] proposed a distributional version of off-policy Q($\lambda$) and TD($\lambda$) algorithms, and Wu et al. [68] proposed a distributional version of fitted Q evaluation to solve the distributional offline policy evaluation problem. Wiltzer et al. [67] proposed an approach for evaluating the return distributions for all policies simultaneously when the reward is deterministic or in the finite-horizon setting. Wiltzer et al. [66] studied distributional policy evaluation in the multivariate reward setting and proposed corresponding TD learning algorithms. Beyond the tabular setting, Bellemare et al. [3], Lyle et al. [32], Bellemare et al. [4] proposed various distributional TD learning algorithms with linear function approximation under different parametrizations.
A series of recent studies have focused on the theoretical properties of distributional TD learning. Rowland et al. [48], Böck and Heitzinger [7], Zhang et al. [73], Rowland et al. [50, 51], Peng et al. [43] analyzed the asymptotic and non-asymptotic convergence of distributional TD learning (or its model-based variants) in the tabular setting. Among these works, Rowland et al. [51], Peng et al. [43] established that in the tabular setting, learning the full return distribution is statistically as easy as learning its expectation in the model-based and model-free settings, respectively. And Bellemare et al. [3] provided an asymptotic convergence result for categorical TD learning with linear function approximation.
Beyond the problem of distributional policy evaluation, Rowland et al. [49], Wang et al. [63, 64] showed that theoretically the classic value-based reinforcement learning could benefit from distributional reinforcement learning. Bäuerle and Ott [1], Chow and Ghavamzadeh [11], Marthe et al. [33], Noorani et al. [40], Moghimi and Ku [35], Pires et al. [44] considered optimizing statistical functionals of the return, and proposed algorithms to solve this harder problem.
Stochastic Approximation.
Our Linear-CTD falls into the category of LSA. The classic TD learning, as one of the most classic LSA problems, has been extensively studied [5, 61, 6, 14, 41, 15, 25, 26, 53, 69]. Among these works, Li et al. [26], Samsonov et al. [54] provided the tightest bounds for Linear-TD with constant step sizes, which is also considered in our paper. While Wu et al. [69] established the tightest bounds for Linear-TD with polynomial-decaying step sizes.
For general stochastic approximation problems, extensive works [23, 56, 37, 38, 20, 28, 16, 54, 10] have provided solid theoretical understandings.
Appendix C Omitted Results and Proofs in Section 2
C.1 Linear Projected Bellman Equation
@@
$=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s)\left({{\mathbb{E}}\left[{r_{0}\mid s_{0}=s}\right]+\gamma{\mathbb{E}}\left[{{\bm{\phi}}(s_{1})^{\top}\mid s_{0}=s}\right]{\bm{\psi}}}\right)}\right]$
$=$ ${\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s,s^{\prime}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right]{\bm{\psi}}+{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s,r}\left[{{\bm{\phi}}(s)r}\right],$
@@
$\left({{\bm{\Sigma}}_{{\bm{\phi}}}-\gamma{\mathbb{E}}_{s,s^{\prime}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right]}\right){\bm{\psi}}={\mathbb{E}}_{s,r}\left[{{\bm{\phi}}(s)r}\right].$
C.2 Convergence Results for Linear TD Learning
@@
In Linear-TD, our goal is to find a good estimator $\hat{{\bm{\psi}}}$ such that $\left\|{{\bm{V}}_{\hat{{\bm{\psi}}}}-{\bm{V}}_{{\bm{\psi}}^{\star}}}\right\|_{\mu_{\pi}}=\left\|{\hat{{\bm{\psi}}}-{\bm{\psi}}^{\star}}\right\|_{{\bm{\Sigma}}_{{\bm{\phi}}}}\leq\varepsilon$. [54] considered the Polyak-Ruppert tail averaging $\bar{{\bm{\psi}}}_{T}:=\left({T/2+1}\right)^{-1}\sum_{t=T/2}^{T}{\bm{\psi}}_{t}$, and showed that in the generative model setting with constant step size $\alpha\simeq(1-\gamma)\lambda_{\min}$,
$T=\widetilde{{\mathcal{O}}}\left({\frac{\left\|{{\bm{\psi}}^{\star}}\right\|^{2}_{{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)^{2}\lambda_{\min}}\left({\frac{1}{\varepsilon^{2}}+\frac{1}{\lambda_{\min}}}\right)}\right)$
is sufficient to guarantee that $\left\|{{\bm{V}}_{\bar{{\bm{\psi}}}_{T}}-{\bm{V}}_{{\bm{\psi}}^{\star}}}\right\|_{\mu_{\pi}}\leq\varepsilon$. They also provided sample complexity bounds when taking the instance-independent (i.e., not dependent on unknown quantity) step size, and in the Markovian setting.
C.3 Categorical Parametrization is an Isometry
Proposition C.1.
The affine space $\left({{\mathscr{P}}^{\mathsf{sign}}_{K},\ell_{2}}\right)$ is isometric with $\left({{\mathbb{R}}^{K},\sqrt{\iota_{K}}\left\|{\cdot}\right\|_{{\bm{C}}^{\top}{\bm{C}}}}\right)$, in the sense that, for any $\nu_{{\bm{p}}_{1}},\nu_{{\bm{p}}_{2}}\in{\mathscr{P}}^{\mathsf{sign}}_{K}$, it holds that $\ell_{2}^{2}(\nu_{{\bm{p}}_{1}},\nu_{{\bm{p}}_{2}})=\iota_{K}\left\|{{\bm{p}}_{1}-{\bm{p}}_{2}}\right\|^{2}_{{\bm{C}}^{\top}{\bm{C}}}$, where ${\bm{C}}$ is defined in Eqn. (11).
@@ C.3 Categorical Parametrization is an Isometry
$=$ $\iota_{K}\left\|{{\bm{p}}_{1}-{\bm{p}}_{2}}\right\|^{2}_{{\bm{C}}^{\top}{\bm{C}}}.$
C.4 Categorical Projection Operator is Orthogonal Projection
Proposition C.2.
[4, Lemma 9.17] For any $\nu\in{\mathscr{P}}^{\mathsf{sign}}$ and $\nu_{{\bm{p}}}\in{\mathscr{P}}^{\mathsf{sign}}_{K}$, it holds that
@@ C.4 Categorical Projection Operator is Orthogonal Projection
C.5 Categorical Projected Bellman Operator
@@
${\bm{p}}_{\bm{\mathcal{T}}^{\pi}\bm{\eta}}(s)=$ ${\mathbb{E}}\left[{{\bm{g}}_{K}(r_{0})+\left({{\bm{G}}(r_{0})-\bm{1}_{K}^{\top}\otimes{\bm{g}}_{K}(r_{0})}\right){\bm{p}}_{\bm{\eta}}(s_{1})\Big{|}s_{0}=s}\right]$
@@
This proposition is a special case of Proposition D.3, whose proof can be found in Appendix D.4.
Appendix D Omitted Results and Proofs in Section 3
D.1 Linear-Categorical Parametrization is an Isometry
Proposition D.1.
The affine space $\left({{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K},\ell_{2,\mu_{\pi}}}\right)$ is isometric with $\left({{\mathbb{R}}^{dK},\sqrt{\iota_{K}}\left\|{\cdot}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}}\right)$, in the sense that, for any $\bm{\eta}_{{\bm{\theta}}_{1}},\bm{\eta}_{{\bm{\theta}}_{2}}\in{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}$, it holds that $\ell_{2,\mu_{\pi}}^{2}\left({\bm{\eta}_{{\bm{\theta}}_{1}},\bm{\eta}_{{\bm{\theta}}_{2}}}\right)=\iota_{K}\left\|{{\bm{\theta}}_{1}-{\bm{\theta}}_{2}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}$.